OpenStudy (wesdg1978):

Suppose that f '(x)=2xf(x) and f(2)=5. How do you find g '(pi/3) if g(x)=f(secx)

hartnn (hartnn):

do u know chain rule, (i guess you know) g(x) = f(sec x) g'(x) =...?

OpenStudy (wesdg1978):

I thought chain rule was with something raised to a power?

hartnn (hartnn):

chain rule can be applied to anything of the form \(f(g(x))\) a function within a function, a composite function here f(sec x) is composite function of f(x) and sec x

OpenStudy (wesdg1978):

So do I take the derivative of sec x? If f '(x)=2xf(x), 2x is the derivative of x^2 but does that help with this problem? I know I'm probably over thinking this and making it harder than it is.

hartnn (hartnn):

g(x) = f(sec x) differentiating this w.r.t. x \(g'(x)= [f(sec x)]' = f'(\sec x) \dfrac{d}{dx}\sec x\) got this chain rule part ?

OpenStudy (wesdg1978):

Yes

hartnn (hartnn):

and whats derivative of sec x ?

OpenStudy (wesdg1978):

sec x tan x

hartnn (hartnn):

now to get g'(pi/3), just plug in x = pi/3 ! what u get ?

OpenStudy (wesdg1978):

Let me go back a step for a second,  Is this right?